OpenStudy (anonymous):

Which of the following choices is the variance of the sample shown here? 30, 32, 47, 57, 66, 79 A. 205.5 B. 17.6 C. 19.2 D. 371.8 E. 309.2

Elsa213 (elsa213):

Mean of sample = (30+32+47+57+66+79) / 6 = 51.83 Variance = Σ (mean - each_sample)² divided by (sample size minus 1) = (476.69 + 393.36 + 23.36 + 26.69 + 200.69 + 738.03) / 5 = 371.8 (D) is the correct choice The above is technically called the "sample variance"; there is also a "population variance" where we would divide by the sample size instead of 1 less. In that case, the result would be 309.8, but the sample variance is the more frequently used quantity (It's also closer to the corresponding choice above)

Elsa213 (elsa213):

@countryrockbabe

OpenStudy (anonymous):

Yawn

OpenStudy (anonymous):

thanks

Elsa213 (elsa213):

May I have medal?

OpenStudy (anonymous):

how do i do that?

Elsa213 (elsa213):

Click on best response :)

Elsa213 (elsa213):

Thank you :D

OpenStudy (anonymous):

oh lol haha i got it (: cool thanks im new at ths

Elsa213 (elsa213):

Thats ok. :) the Same thing happened 2 me when I was new :D

OpenStudy (anonymous):

Yawn